OpenStudy (pphalke):

Nico is saving money for his college education. He invests some money at 6% and $1900 less than that amount at 5%. the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate?

OpenStudy (hiimcasey):

What is the middle number when the numbers {8, –2.4, 3.2, –1.5, 1.3} are ordered from least to greatest?

OpenStudy (pphalke):

@hiimcasey ask this in the ask question box plz

OpenStudy (pphalke):

@Janu16 help plz

OpenStudy (janu16):

ok hold plz

OpenStudy (janu16):

.06x+.05(x-1,900)=191 .11x-95=191 .11x=191+95 .11x=286 x=2600 invested at 6% 2600-1900=700 amount invested at 5% so he invested 2600 at 6% and invested 700 st 5%

OpenStudy (janu16):

@pphalke

OpenStudy (pphalke):

thank u